40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 16 21:13:07 2012, Suresh, Update, Electronics, REFL165 repair: PD replaced, DC response checked with a torch light PC_30641_old.jpgPD_30641_new.jpgIMG_0548.JPG
    Reply  Tue Feb 28 16:31:08 2012, Suresh, Update, Electronics, REFL165 repair: Characterization REFL165_Characterization.pdfREFL165_response_shift.pdf
       Reply  Wed Feb 29 00:22:35 2012, Suresh, Update, Electronics, REFL165 repair: Installed on the AS table 
       Reply  Wed Feb 29 01:14:40 2012, Suresh, Update, Electronics, REFL165 repair: Characterization 
Message ID: 6290     Entry time: Thu Feb 16 21:13:07 2012     Reply to this: 6333
Author: Suresh 
Type: Update 
Category: Electronics 
Subject: REFL165 repair: PD replaced, DC response checked with a torch light 

[Koji, Suresh]

Kiwamu mentioned that REFL165 is not responding and its DC out seems saturated at 9V.  Koji and I checked to see if changing the power supply to the PD changed its behaviour. It did not.  

I then look a close look at the PD and found that the front window of the PD was not clear and transparent.  There was a liquid condensation inside the window, indicating an over heating of the PD at some point.  It could have arisen due to excessive incident power.  The pic below shows this condensation:

PC_30641_old.jpg

 

I also checked the current flowing through the reverse bias voltage line.  There was a voltage drop of 3V across R22 (DCC D980454-01-C)   indicating a 150mA of current through the PD.  This is way too much above the operating current of about 20mA.   The diode must have over heated.

I pulled out the old PD out and installed a new one from stock.  The pic below shows the clear window of a new PD.

PD_30641_new.jpg

After changing the PD I checked the DC output voltage while shining a torch light on to the PD.  It showed an output of about 30 to 40 mV.  This seemed okay because the larger 2mm photodiodes showed ~100mA DC output with the same torch.Below is the current state of the ckt board.

IMG_0548.JPG

 

I will tune the PD to 165 MHz tomorrow and measure its transimpedance.

ELOG V3.1.3-